OpenStudy (anonymous):

y=tan^-1(5ax/a^2-6x^2)show that dy/dx=3a/a^2+9x^2+2a/a^2+4x^2

OpenStudy (chaise):

If you write this in equation editor I might be able to help.

OpenStudy (anonymous):

can you write the equation more simple???

OpenStudy (jhannybean):

@hartnn

OpenStudy (jhannybean):

quotient rule? I have no idea why it's dy/dx....

hartnn (hartnn):

my first thoughts will be tan y =(5ax/(a^2-6x^2)) ....find sec^2 y from here, then differentiate both sides.... i'll try on paper first whether this works....

OpenStudy (jhannybean):

I have been...i don't understand where dy comes into the picture...lol

hartnn (hartnn):

when you differentiate y, with respect to x, you get dy/dx

OpenStudy (jhannybean):

so you're solving for y '...

hartnn (hartnn):

since, they have written dy/dx and not, \(\partial y/\partial x\), we need to assume that a is constant here..

hartnn (hartnn):

yes, we are proving that by solving for y'

OpenStudy (jhannybean):

\[\large \tan^{-1}(\frac{5ax}{a^2-6x^2})\]Show that \[\large \frac{dy}{dx}=\frac{3a}{a^2}+9x^2+\frac{2a}{a^2}+4x^2\]

OpenStudy (jhannybean):

Omg this is too lengthy.

OpenStudy (nincompoop):

you signed up for it :P

OpenStudy (nincompoop):

I think it is supposed to be [3a/(a^2+9x^2)]+[2a/(a^2+4x^2)]

OpenStudy (nincompoop):

I might be wrong since I don't know this stuff :X